OpenStudy (anonymous):

Which is a completely factored form of 27x^3+y^6 1. (3x+y^2)^3 2. (3x+y^2)(9x^2+y^4) 3. (3x-y^2)(9x^2+3xy^2+y^4) 4. (3x+y^2)(9x^2-3xy^2+y^4)

OpenStudy (anonymous):

can u tell me what does \(a^3 + b^3 \) equal to? u know \(27x^3 + y^6 = (3x)^3 + (y^2)^3 \)

OpenStudy (anonymous):

I have no idea how to do this Im trying to learn

OpenStudy (anonymous):

well,\[x^3 + b^3 = (a+b)(a^2+b^2 - ab )\]

OpenStudy (anonymous):

now put \(a=3x \) and \(b=y^2 \) in that equation

OpenStudy (anonymous):

* sry \(a^3\) i put \(x^3\) by mistake

OpenStudy (anonymous):

ok but where does the y^2 come from

OpenStudy (anonymous):

\(y^6 = (y^2)^3 \)

OpenStudy (anonymous):

oh ok

Nnesha (nnesha):

take cube root for fist one and then take cube of second

Nnesha (nnesha):

wait  just write here don't worry about the signs yet i will tell you that later